OpenStudy (anonymous):

which of the following points do not lie on the circle whose equation is (x-3)^2+(y-2)^2=100

OpenStudy (accessdenied):

i think if you test the points, you'll be able to figure it out (because if the equation isn't equal for the given point, then the equation isn't defined for that point)

Directrix (directrix):

Where is the list of following points?

OpenStudy (anonymous):

1. (-7,2) 2. (3,-8) 3. 3,12) 4. (13,1)

OpenStudy (anonymous):

can you show me how to do this problem if possible?

Directrix (directrix):

I tested points. My first test was on (-3,8). (x-3)^2+(y-2)^2=100 ( (-3) - 3)^2 + ( 8 - 2) ^ 2 == 100 ? Yes or no?

Directrix (directrix):

0^ + 6^2 = 100? NO. The point (-3,8) is not on the graph.

OpenStudy (anonymous):

Substitute the x and y values from the ordered pair into the equation, then see if it's true.

Directrix (directrix):

Take a look at the graph at http://www.wolframalpha.com/input/?i=%28x-3%29%5E2%2B%28y-2%29%5E2%3D100

OpenStudy (anonymous):

thank you, I got 4. (13,1)

Directrix (directrix):

Any time.